• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name LeRose

    Log entry time 16:01:55 on June22,2000

    Entry number 44672

    keyword=Shift Summary

    Plan is to do production running at thetacm=-50 degrees
    Electron: P0=3.660 @14.075 deg collimator open
    Hadron: P0=1.27 @50.179 deg collimator open
    15cm H2 target
    Begin with run run 2446 in progress Stopped at 8:20 because accelerator having RF problems (0.69M events)
    8:25 Started 2447
    9:08 Ended 2447 (1.3M). 2448 started
    9:30 ended 2448 downstream 3x problem traced to loose cable. Reconnected cable.
    9:38 ended 2449 short run to test 3x downstream problem. Repaired!
    10:20 ended 2450 (1.00M)
    Accelerator doing some steering, going to pulsed beam for a few minutes. Data taking suspended.
    10:34 beam back
    11:20 More beam problems. Vertical moton in the 0L0 region. Estimate 15 minutes down, replacing trim cards.
    12:00 HAC rebooted (was loaded down with excessive tasks)
    12:05 beam back, 2452 started.
    12:58 2452 ends. Beam is still not very stable (on a minute, off a half minute, on three minutes, off again etc....) Operations elog (114768) says vertical motion is still there.
    13:17 MCC has an IOC problem (IOCea5) with a magnet estimate 10 minutes down. Ended run 2453.
    13:35 IOC went down again. Estimate another 10-15 minutes down.
    14:15 Still waiting
    14:30 MCC called. They are having multiple problems. Those problems conflict with the planned spot move (scheduled for 1400) and are holding it up. When the problems are resolved they will make the spot move. No predictions for when we'll see beam again.


    2446...0.69M
    2447...1.30M
    2448...0.50M
    2449...0.15M
    2450...1.00M
    2451...0.80M
    2452...1.30M
    2453...0.44M

    1.4 Coulombs up to now